OpenStudy (anonymous):

3x-2y = 11 6x + 11y = 97 solve using the elimination method .Show your work .if the system has no solution or an infinite number of solution ,state this?

OpenStudy (anonymous):

3x-2y=11 => 6x-4y=22 => -6x+4y=-22 6x+11y=97 Add the two. we get, 15y=75 =>y=5 3x-2*5=11 3x=21 x=7

OpenStudy (anonymous):

multiply first equation by 2 subtrct to eliminate x solve for y plug va,ue of y into one equn. to get x

OpenStudy (anonymous):

\[6x+2y+3+11y=\]